A 75 kg skydiver can be modeled as a rectangular "box" with dimensions 24 cm × 41 cm × 170 cm . If he falls feet first, his drag coefficient is 0.80.

Part A What is his terminal speed if he falls feet first? Use ρ = 1.2 kg/m3 for the density of air at room temperature. Express your answer with the appropriate units.
Answer #1

We know that Drag force is

D = 1/2*density*v2*C*A

Where C is the Drag coeffecient
W = m*g

At terminal speed, Fnet = 0,

D - W = 0,

D = W

So, solving for V

mg=0.5d*v2*C*A

v=(2mg/CAd)0.5
A is the cross sectional area going into the wind
S = .24 m * .41 m
C = 0.80

So,

v=(2*75*9.81/0.8*0.24*0.41*1.2)0.5

v=124.81 m/s